# Schrodinger’s Memory: Large Language Models

### v1

@@ Abstract
Memory is the foundation of LLMs’ functionality, yet past research has lacked an in-depth exploration of their memory capabilities and underlying theory. In this paper, we apply UAT theory to explain the memory mechanism of LLMs and propose a new approach for evaluating LLM performance by comparing the memory capacities of different models. Through extensive experiments, we validate our theory and the memory abilities of LLMs. Finally, we compare the capabilities of the human brain and LLMs, highlighting both their similarities and differences in terms of working mechanisms.
@@ 1 Introduction
Language is not only one of humanity’s most essential capabilities but also the foundation of communication, knowledge transfer, and the development of civilization. Language models can be seen as a simulation of human intelligence, enabling them to perform tasks that traditionally required human input. Large Language Models (LLMs), particularly those based on the Transformer architecture, are among the hottest research topics in artificial intelligence today. They have begun to mimic human language abilities and are already impacting various aspects of everyday life, such as machine translation, text summarization, sentiment analysis, question-answering systems, and text generation.
@@
Despite the impressive capabilities of LLMs, the research on their memory mechanisms remains underexplored. Memory is a crucial ability for humans; without it, we would struggle to complete even the simplest tasks. For instance, something as routine as eating requires remembering where to eat, how to get there, what to eat, and which utensils to use. These everyday memories have become second nature to us, so much so that we rarely think about them consciously. Clearly, memory plays a vital role in guiding nearly every aspect of our daily lives.
@@
As LLMs become more powerful, it raises the question: do these models possess memory? If so, what form does it take? How does it differ from human memory? Current research into LLM memory mechanisms primarily follows two directions:
@@
Expanding Context Length: This approach aims to equip LLMs with long-term memory by extending the context window Chen et al. (2023); Zhu et al. (2023); Yang (2023); Fei et al. (2023). Since short contexts fail to provide enough information, increasing the context length allows the model to maintain more comprehensive information across long sequences.
@@
External Memory Integration: This method involves building memory storage systems Graves et al. (2014); Xiao et al. (2024); Wu et al. (2022); Yang et al. (2024b) that encode and store past events Zhang et al. (2023), allowing the model to retrieve and update memories as needed. Such mechanisms enable models to forget or reinforce certain memories over time.
@@
While these studies have made strides in addressing the memory limitations of LLMs, they do not fully explain how memory works within these models. For example, when asked ”Who is the President of the United States?”, LLMs like GPT-4 or Llama-3 might answer with outdated information, indicating that some form of memory is present. However, this memory isn’t from an external storage unit but is inferred from the model’s training data. This raises fundamental questions: Why do LLMs exhibit this inferential memory? How does it compare to human memory?
This paper will leverage the Universal Approximation Theorem (UAT) to explain the reasoning and memory abilities of LLMs. The contributions of this work are as follows:
  • An explanation of LLMs’ memory and reasoning abilities through the lens of UAT.
@@
  • A new, more objective method for evaluating LLMs’ memory and reasoning strengths.
@@
  • A comparison between the memory and reasoning capabilities of LLMs and human cognition.
@@ 2 UAT and LLMs
The UAT (Cybenko, 2007; Popescu et al., 2009) serves as the foundational theory of deep learning. Our goal is to theoretically explain Transformer-based LLMs using the UAT framework. To do this, we will first present the mathematical form of UAT in Section 2.1, followed by the corresponding UAT form for LLMs in Section 2.2. We will then use this UAT form to explain the memory characteristics of LLMs.
@@ 2.1 UAT
In this section, we will briefly introduce the UAT. The UAT was originally introduced by Cybenko (2007). According to Theorem 2 from Cybenko (2007), if $\sigma$ is any continuous sigmoidal function, then a finite sum of the following form:
@@
This indicates that, with a sufficiently large $N$, a neural network can approximate any continuous function over a closed interval. Hornik et al. (1989) further establishes that multilayer feedforward networks adhere to the UAT, capable of approximating arbitrary Borel measurable functions. Considering Equation (1), where the function $G(\mathbf{x})$ outputs a scalar in $\mathbb{R}$, the framework naturally extends when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation across each dimension. To handle this multidimensional output, a straightforward modification to Equation (1) is sufficient: the transformation matrix $\mathbf{W}_{j}$ is adjusted to lie in the space $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is redefined as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
@@ 2.2 The UAT Format of Transformer-Based LLMs
Figure 1: The process of Transformer.
Current LLMs are primarily based on the memory-enhanced Transformer architecture. In UAT2LLMs Wang & Li (2024b), it has already been demonstrated that the mathematical structure of multi-layer Transformers aligns with the UAT in a general sense. However, unlike the original UAT, the UAT form of Transformer-based models has the ability to dynamically fit functions based on the input. Figure 1 illustrates a basic Transformer module, and according to UAT2LLMs, the corresponding UAT form for Figure 1 is:
@@
where $\mathbf{x}_{i+1}$ is the output of ${i+1}$-th layer, $\mathbf{x}_{0}$ is the input of the network, $\mathbf{b}^{\prime}_{j,2}=(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$. The term $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$ layer of UAT with $\mathbf{x}_{0}$ as input. This enhances the model’s ability to dynamically adjust functions based on input. In the multi-head attention mechanism, the parameters change dynamically with the input. Therefore, in the formula above, all $\mathbf{W}^{\prime}_{j,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ parameters in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
@@ 3.1 The Definition of Memory
Before formally studying memory in LLMs, it’s important to first define or provide a relatively accurate description of memory. According to Wikipedia:
@@
However, this definition has some fundamental issues. Encoding data or information is not problematic, as information in the brain is transmitted via electrical signals, and we need to encode that information in a way the brain can process. The problem arises with the concepts of ”storage” and ”retrieval.” The brain does not have a structure analogous to a database for storing information. So, where is this information actually stored? Is it in the neurons of the brain? If so, does a single neuron store a word, or does it store an entire sentence?
@@
Therefore, we need a more precise definition of the concept of ”memory”:
@@
  • Input: The input is information that is the same or similar to what the brain (or an LLM) has previously encountered (this is a necessary condition for memory—without input, there is no memory).
  • Output: The result based on the input, which could be correct, incorrect, or forgotten. If the result is correct, it means it aligns with information previously acquired. (Note: for there to be an output, there must be an input; memory doesn’t emerge spontaneously. Without specific input conditions, a person wouldn’t recall a particular event.)
@@ 3.2 The Memory Mechanism and Ability of LLMs
In Section 2.2, we have introduced the UAT format corresponding to Transformer-based LLMs. This UAT format can dynamically adjust to fit the corresponding output based on the input. Following this idea, could we also consider that the memory feature of LLMs adjusts to specific outputs based on the input? To explore this hypothesis, we designed a simple experiment.
@@
To test the memory ability of LLMs, we fine-tuned the CN Poems and ENG Poems on Qwen series models Bai et al. (2023); Yang et al. (2024a) and bloom Workshop et al. (2023) models. The results in Table 1 indicate that LLMs possess memory capabilities and align perfectly with the definition of memory we provided. The training process is akin to giving a person 2,000 poems and asking them to memorize as many as possible, with a limit of reading each poem only 100 times. In the CN Poems dataset, the best-performing models were Qwen2-1.5B-Instruct and bloom-1b4-zh, which remembered 1,938 and 1,932 poems respectively. In contrast, for the ENG Poems dataset, nearly all models memorized all the poems.
These results are remarkable. A human, without specialized memory training, would struggle to memorize even 100 poems under such conditions, whereas the LLMs were able to memorize almost 100% of the 2,000 poems. However, overall performance on the CN Poems dataset was weaker for models like Qwen1.5-0.5B-Chat, Qwen2-0.5B-Instruct, and bloom-389m-zh. We believe this is due to insufficient pre-training, resulting in relatively poor language comprehension. For instance, Qwen2-0.5B-Instruct outperformed Qwen1.5-0.5B-Chat, even though both models had similar data and model sizes.
Stronger language understanding suggests a better fit for natural language, which can be interpreted as a model having a higher language comprehension ability. This, in turn, aids in model expansion. The original Qwen documentation supports this, as Qwen2-0.5B-Instruct outperforms Qwen1.5-0.5B-Chat across a variety of tasks. Similarly, a person with better language comprehension would learn other related information more efficiently. The availability and comprehensiveness of English-language data compared to Chinese likely contributed to the models’ superior performance in fitting English texts.
@@
Figure 2 shows the results of generating a full poem based on partial input information. As seen in the figure, after training, the models can successfully recite the entire poem using only the title and author. Figure 2 displays some examples of prediction errors.
Figure 2: The process of Transformer.
Figure 3: The process of Transformer.
@@ 4 A Comparision Between Human and LLMs
Reasoning Ability: The ability to generate results consistent with previously learned knowledge based on specific inputs.
From this definition, we can identify several key factors in reasoning ability: the knowledge learned, the specific input, and the ability to produce results aligned with the learned knowledge. According to this definition, the memory capabilities of LLMs can also be considered a form of reasoning. As shown in Figure 2, LLMs can iteratively learn to reconstruct a complete poem based solely on its title and author. These poems are not stored in a particular weight within the model but are dynamically generated based on the input. This is why we describe LLMs’ memory as ”Schrödinger’s memory”—we can only determine if an LLM remembers something after posing a question and examining the output. Otherwise, it is impossible to know. Humans operate similarly: we can only verify our memory by answering specific questions; otherwise, we cannot assess it. For example, if you ask someone how many poems they remember, they might not be able to give a precise answer, but if you ask them if they remember a specific poem, they can usually respond.
As shown in Figure 3, although some predictions are incorrect, the results still align with linguistic conventions and match the poem’s title to some degree. This could be seen as a form of creativity in the absence of exact memory, which is also a form of reasoning.
@@
  • Model Size: In practice, larger LLMs tend to be more powerful. Theoretically, as shown in the UAT2LLMs Wang & Li (2024b) and UAT2Parallel Wang & Li (2024a) models, the more layers a model has, the stronger its dynamic fitting ability becomes, resulting in better performance.
@@
A great example of dynamic fitting in the brain is Henry Molaison Scoville & Milner (1957); Victor et al. (1961); Milner & Klein (2015). After his hippocampus Bliss & Collingridge (1993); Squire (1992); Erickson et al. (2011); Eckardt (1980) was damaged, he could no longer form new long-term memories, though his existing memories remained intact. We believe that the hippocampus acts as a switch controlling whether the weights responsible for long-term memory in the brain can be updated. Once the hippocampus is damaged, the brain’s weight parameters can no longer change, meaning that while past inputs (before the hippocampal damage) can still produce corresponding outputs (i.e., recalling past events), the inability to update weights prevents the formation of new memories.

### v2

@@ Abstract
Memory is the foundation of all human activities; without memory, it would be nearly impossible for people to perform any task in daily life. With the development of Large Language Models (LLMs), their language capabilities are becoming increasingly comparable to those of humans. But do LLMs have memory? Based on current performance, LLMs do appear to exhibit memory. So, what is the underlying mechanism of this memory? Previous research has lacked a deep exploration of LLMs’ memory capabilities and the underlying theory. In this paper, we use Universal Approximation Theorem (UAT) to explain the memory mechanism in LLMs. We also conduct experiments to verify the memory capabilities of various LLMs, proposing a new method to assess their abilities based on these memory ability. We argue that LLM memory operates like Schrödinger’s memory, meaning that it only becomes observable when a specific memory is queried. We can only determine if the model retains a memory based on its output in response to the query; otherwise, it remains indeterminate. Finally, we expand on this concept by comparing the memory capabilities of the human brain and LLMs, highlighting the similarities and differences in their operational mechanisms.
@@ 1 Introduction
Language is not only one of humanity’s most important abilities but also the foundation of communication, knowledge transfer, and the development of civilization. Language models can be seen as simulations of human intelligence, enabling them to perform tasks traditionally achievable only by humans. In particular, LLMs based on the Transformer architecture have become one of the hottest topics in artificial intelligence research today. These models have acquired some human-like language capabilities and are already impacting daily life in areas such as machine translation, text summarization, sentiment analysis, question-answering systems, and text generation.
@@
Despite their impressive performance, the study of LLMs’ memory mechanisms remains insufficient. Memory is a critical ability for humans; without it, we wouldn’t be able to complete even the simplest tasks. For example, routine activities like eating require us to remember where to eat, how to get there, what to eat, and which utensils to use. These everyday memories have become second nature to the extent that we rarely think about them consciously. Clearly, memory plays a crucial role in guiding various aspects of our daily lives.
@@
As LLMs become increasingly powerful, this raises an important question: Do these models possess memory capabilities? If so, in what form? How does it differ from human memory? Current research on LLM memory mechanisms generally focuses on two main directions:
@@
Expanding Context Length: This approach aims to equip LLMs with long-term memory by extending the context window (Chen et al., 2023; Zhu et al., 2023; Yang, 2023; Fei et al., 2023). Since short contexts fail to provide enough information, increasing the context length allows the model to maintain more comprehensive information across long sequences.
@@
External Memory Integration: This method involves building memory storage systems (Graves et al., 2014; Xiao et al., 2024; Wu et al., 2022; Yang et al., 2024b) that encode and store past events (Zhang et al., 2023), allowing the model to retrieve and update memories on disks as needed. Such mechanisms enable models to forget or reinforce certain memories over time.
@@
Although these studies have made progress in addressing the memory limitations of LLMs, they have not fully explained how memory functions within these models. For example, when asked, ”Who is the President of the United States?” LLMs like GPT-4 (openai2024gpt4technicalreport) or Llama-3 (dubey2024llama3herdmodels) may sometimes provide outdated information, indicating that some form of memory is indeed present. However, this memory does not come from an external storage unit but is inferred by the model based on the input. This raises fundamental questions: Why do LLMs exhibit this ability to infer previously learned information from the input? How does this differ from human memory?
In this paper, we use UAT theory to explain this ability of recalling information learned from the past based on input cues. We argue that this information can be understood as a dynamic approximation capability of UAT, where the model fits the input to a corresponding result, and the observed phenomenon appears as memory. We call this ”Schrödinger’s memory” because we can only determine whether the LLM has this memory by asking it and analyzing its response; otherwise, the memory remains indeterminate. Additionally, we evaluate the memory capabilities of several models and propose that this approach can be used to assess the overall ability of LLMs. The contributions of this work are as follows:
  • We explain LLMs’ memory and reasoning abilities through the lens of UAT.
@@
  • We propoese a new, objective method for evaluating LLMs’ capabilities: memory ability assessment.
@@
  • We do a comparison between the memory and reasoning capabilities of LLMs and human cognition.
@@ 2 UAT and LLMs
The UAT (Cybenko, 2007; Popescu et al., 2009) serves as the foundational theory of deep learning. Our goal is to theoretically explain memory of Transformer-based LLMs using the UAT framework. To do this, we will first present the mathematical form of UAT in Section 2.1, followed by the corresponding UAT form for LLMs in Section 2.2. We will then use this UAT form to explain the memory abilities of LLMs.
@@ 2.1 UAT
In this section, we provide a brief overview of the UAT, which was first proposed by Cybenko (2007). As stated in Theorem 2 by Cybenko (2007), if $\sigma$ represents any continuous sigmoidal function, then a finite sum of the following form:
@@
This suggests that, for a sufficiently large $N$, a neural network can approximate any continuous function on a closed interval. Hornik et al. (1989) further demonstrates that multilayer feedforward networks conform to the UAT, enabling them to approximate arbitrary Borel measurable functions. In the context of Equation (1), where the function $G(\mathbf{x})$ produces a scalar in $\mathbb{R}$, this framework naturally generalizes when $G(\mathbf{x})$ maps to $\mathbb{R}^{m}$, requiring approximation for each dimension. To accommodate this multidimensional output, a simple adjustment to Equation (1) is needed: the transformation matrix $\mathbf{W}_{j}$ is modified to reside in $\mathbb{R}^{n\times m}$, the bias term $\theta_{j}$ is redefined as a vector in $\mathbb{R}^{m}$, and $\alpha_{j}$ is reshaped into a matrix.
@@ 2.2 The UAT Format of Transformer-Based LLMs
Current LLMs are primarily based on Transformer architecture. In UAT2LLMs (Wang & Li, 2024b), it has already been demonstrated that the mathematical structure of multi-layer Transformers aligns with the UAT in a general sense. However, unlike the original UAT, the UAT form of Transformer-based models has the ability to dynamically fit functions based on the input. Figure 1 illustrates a basic module in Transformer, and according to UAT2LLMs, the corresponding UAT form for Figure 1 is:
@@
where $\mathbf{x}_{i+1}$ represents the output of the ${i+1}$-th layer, with $\mathbf{x}_{0}$ as the network’s input. The term $\mathbf{b}^{\prime}_{j,2}$ is computed as $(\mathbf{W}^{\prime}_{j,2}\mathbf{b}^{\prime}_{j-1,3}+\mathbf{b}^{\prime}_{j,2})+\mathbf{W}^{\prime}_{j,2}UAT_{j-1}^{R}$, where $UAT_{j-1}^{R}=\sum_{k=1}^{j-1}\mathbf{W}^{\prime}_{k,3}\sigma(\mathbf{W}^{\prime}_{k,2}\mathbf{x}^{\prime}_{0}+\mathbf{b}^{\prime}_{k,2})$. The value of $\mathbf{b}^{\prime}_{j,2}$ is approximated by the $j$-th layer of the UAT, with $\mathbf{x}_{0}$ as the input. This allows the model to dynamically adjust functions based on the input. In the multi-head attention mechanism, the parameters are modified dynamically in response to the input. Therefore, in the formula above, all parameters $\mathbf{W}^{\prime}_{j,1}$, $\mathbf{W}^{\prime}_{j,2}$, and $\mathbf{W}^{\prime}_{j,3}$ in layer $i$, where $j=1,\ldots,i$, are dynamically adjusted based on the input.
Figure 1: The basic block in Transformer-based LLMs.
@@ 3.1 The Definition of Memory
Before delving into the study of memory in LLMs, it is important to first define or provide a relatively precise description of what memory is. According to Wikipedia:
@@
However, this definition has some fundamental issues. Encoding data or information is not problematic, as information in the brain is transmitted via electrical signals, and we need to encode that information in a way the brain can process. The problem arises with the concepts of ”storage” and ”retrieval.” The brain does not have a structure analogous to a database for storing information. So, where is this information actually stored? Is it in the neurons of the brain? If so, does a single neuron store a word, or does it store an entire sentence? Next, we give an exmple:
@@
Therefore, we provide a more precise definition of the concept of ”memory”:
@@
  • Input: To trigger a memory, the input must be the same or similar to information that the brain (or LLM) has previously encountered.
  • Output: The result based on the input, which could be correct, incorrect, or forgotten. If the result is correct, it means it aligns with information previously acquired.
We need to stress that a key requirement for recalling a memory is the presence of input—without input, there is no memory, as memory is activated by input. Even if the brain or an LLM holds a memory of something, it cannot be determined whether that memory exists unless it is prompted by a specific query. Without specific input conditions, a person wouldn’t recall a particular event.
@@ 3.2 The Memory Mechanism and Ability of LLMs
In Section 2.2, we have introduced the UAT format corresponding to Transformer-based LLMs. This UAT format can dynamically adjust to fit the corresponding output based on the input. Following this idea, could we also consider that the memory of LLMs is to fit specific outputs based on the input? To explore this hypothesis, we designed a simple experiment.
@@
Figure 2: The examples of right predictions of two Qwen1.5-0.5B-Chat models which were fine-tuned separately on CN Poems and ENG Poems and subsequently tested the memory ability on their respective datasets, accurately recited the entire poem based on the input.
To test the memory ability of LLMs, we fine-tuned the CN Poems and ENG Poems on Qwen series models (Bai et al., 2023; Yang et al., 2024a) and bloom (Workshop et al., 2023) models. The results in Table 1 indicate that LLMs possess memory capabilities and align perfectly with the definition of memory we provided. The training process is akin to giving a person 2,000 poems and asking them to memorize as many as possible, with a limit of reading each poem only 100 times. In the CN Poems dataset, the best-performing models were Qwen2-1.5B-Instruct and bloom-1b4-zh, which remembered 1,938 and 1,932 poems respectively. In contrast, for the ENG Poems dataset, nearly all models memorized all the poems.
These results are remarkable. A human, without specialized memory training, would struggle to memorize even 100 poems under such conditions, whereas the LLMs were able to memorize almost 100% of the 2,000 poems. However, overall performance on the CN Poems dataset was weaker for models like Qwen1.5-0.5B-Chat, Qwen2-0.5B-Instruct, and bloom-389m-zh. We believe this is due to insufficient pre-training, resulting in relatively poor language comprehension. For instance, Qwen2-0.5B-Instruct outperformed Qwen1.5-0.5B-Chat, even though both models had similar data and model sizes. Stronger language understanding suggests a better fit for natural language, which can be interpreted as a model having a higher language comprehension ability. This, in turn, aids in model expansion. The performances in Qwen documentation (Bai et al., 2023; Yang et al., 2024a) supports this, as Qwen2-0.5B-Instruct outperforms Qwen1.5-0.5B-Chat across a variety of tasks. Similarly, a person with better language comprehension would learn other related information more efficiently. The availability and comprehensiveness of English-language data compared to Chinese likely contributed to the models’ superior performance in fitting English texts.
@@
Figure 2 shows the results of generating a full poem based on partial input information. As seen in the figure, after training, the models can successfully recite the entire poem using only the title and author. Figure 3 displays some examples of prediction errors. Although the prediction was incorrect, it can be observed that the output still aligns with the poem’s title and basic structure.
Based on these results, we believe that LLMs do indeed possess memory, and their memory mechanism works by fitting the input to a specific output. This is why we refer to LLMs’ memory as ”Schrödinger’s memory”—we can only determine whether the LLM has a particular memory when we ask a question and receive a response; otherwise, it remains indeterminate.
Figure 3: The examples of wrong prediction of Qwen1.5-0.5B-Chat which was fine-tuned separately on CN Poems and ENG Poems.
@@ 4 A Comparision Between Human and LLMs
Reasoning Ability: The ability to generate specific results based on previously learned knowledge and specific inputs. The specific results is consist with the previously learned knowledge.
From this definition, we can identify several key factors in reasoning ability: the knowledge learned, the specific input, and the ability to produce results aligned with the learned knowledge. According to this definition, the memory capabilities of LLMs can also be considered a form of reasoning. As shown in Figure 2, LLMs can iteratively learn to reconstruct a complete poem based solely on its title and author. These poems are not stored in a particular weight within the model but are dynamically generated based on the input. This is why we describe LLMs’ memory as ”Schrödinger’s memory”—we can only determine if an LLM remembers something after posing a question and examining the output. Otherwise, it is impossible to know. Humans operate similarly: we can only verify our memory by answering specific questions; otherwise, we cannot assess it. For example, if you ask someone how many poems they remember, they might not be able to give a precise answer, but if you ask them if they remember a specific poem, they can usually respond. As shown in Figure 3, although some predictions are incorrect, the results still align with linguistic conventions and match the poem’s title to some degree. This could be seen as a form of creativity in the absence of exact memory, which is also a form of reasoning.
@@
  • Model Size: In practice, larger LLMs tend to be more powerful. Theoretically, as shown in the UAT2LLMs (Wang & Li, 2024b) and UAT2Parallel (Wang & Li, 2024a) , the more layers a model has, the stronger its dynamic fitting ability becomes, resulting in better performance.
@@
A great example of dynamic fitting in the brain is Henry Molaison (Scoville & Milner, 1957; Victor et al., 1961; Milner & Klein, 2015). After his hippocampus (Bliss & Collingridge, 1993; Squire, 1992; Erickson et al., 2011; Eckardt, 1980) was damaged, he could no longer form new long-term memories, though his existing memories remained intact. We believe that the hippocampus acts as a switch controlling whether the weights responsible for long-term memory in the brain can be updated. Once the hippocampus is damaged, the brain’s weight parameters can no longer change, meaning that while past inputs (before the hippocampal damage) can still produce corresponding outputs (i.e., recalling past events), the inability to update weights prevents the formation of new memories.